OpenStudy (anonymous):

can someone please thoroughly explain each step to this calc problem pleasE??

OpenStudy (jhannybean):

Are you familiar with Reimann Sums?

OpenStudy (anonymous):

yes kinda, i dont have my textbook with me thats why im so stuck

OpenStudy (jhannybean):

\[\int_a^b f(x)dx = \sum_{i=1}^n f(x_i)dx \]

OpenStudy (anonymous):

i just put up an example so someone could explain it step by step and the work so i can apply it to my homework questions(:

OpenStudy (jhannybean):

Start by finding your \(\Delta x\)\[\Delta x = \frac{b-a}{n}\]This will tell you where your right endpoint starts.

OpenStudy (jhannybean):

And the change in your x-value to your next endpoint.

OpenStudy (anonymous):

what is b and a ?

OpenStudy (jhannybean):

Your equation: \(\int_{-2}^2 (x^3+8)dx\) Compare to: \(\int_a^b f(x)dx\)

OpenStudy (anonymous):

1

OpenStudy (jhannybean):

So: \[\Delta x = \frac{b-a}{n} \implies \frac{2-(-2)}{4} = 1\]

OpenStudy (jhannybean):

Yep. so your first rectangle starts at x=1.

OpenStudy (jhannybean):

Since \(\Delta x = 1\), your second rectangle is at x=2

OpenStudy (jhannybean):

Ok, once you find all your right endpoints, you will use the formula \[R_4 = \sum_{i=1}^4 f(x_i)\Delta x\]

OpenStudy (jhannybean):

That means \[\begin{align} R_4 = \sum_{i=1}^4 f(x_i)\Delta x &= \\&= f(x_1)\Delta x +f(x_2)\Delta x +f(x_3)\Delta x + f(x_4)\Delta x \end{align}\]

OpenStudy (anonymous):

so first rectangle is at 1,second is at 2, third is at 3, and fourth is at 4?

OpenStudy (jhannybean):

Yep.

OpenStudy (jhannybean):

Only because the change is \(\Delta x=1\)

OpenStudy (anonymous):

oh this isnt that bad, i get it so far(:

OpenStudy (jhannybean):

Yay!

OpenStudy (jhannybean):

Glad you're comprehending what I am explaining to you.

OpenStudy (jhannybean):

so then for each \(f(x_i)\) you will pluf the \(x\)-value into your original equation: \(x^3+8\) to find the exact value, then multiply it by \(\Delta x\).

OpenStudy (jhannybean):

then add them all up and you're done! :)

OpenStudy (anonymous):

wait i got kinda confused on the last step, can you show me an example?

OpenStudy (anonymous):

or do it with me please??

OpenStudy (jhannybean):

\[x=1~,~2~,~3~,~4\]so find : \(f(1)~,~f(2)~,~f(3)~,~f(4)\)

OpenStudy (jhannybean):

Yes you are right :)

OpenStudy (anonymous):

isnt there two questions? the height and to approximate?

OpenStudy (jhannybean):

No, the total area is 131, the height of each rectangle is just the value you get when you plug in your x value.

OpenStudy (jhannybean):

So technically we answered two questions for the price of one :P

OpenStudy (jhannybean):

Yes :D

OpenStudy (jhannybean):

mmhmm

OpenStudy (anonymous):

ok just making sure, thanks(:

OpenStudy (jhannybean):

And you know why it's an approximate? Because you're only finding the area w.r.t the right endpoints. I'm sure you learned in class that to find the exact area you have to subtract the overestimate by the underestimate?

OpenStudy (anonymous):

yesss

OpenStudy (jhannybean):

That's why :)

OpenStudy (anonymous):

thank you so much, you explained it very well

OpenStudy (jhannybean):

Thank you!